The temperature of a 3.50 L sample of gas is lowered from 355K to 225K.  What will be the resulting volume of the gas?


1
Expert's answer
2021-03-31T08:00:35-0400

"\frac {V1}{T1}= \frac {V2}{T2}"


The resulting volume of gas will be V2


"V2= \frac {V1T2}{T1}"


"V2 = \frac{3.50×225}{355}=2.22L"